The perimeter of a rectangle is 160 cm. The length is 4 less than three times the width. Find the dimensions.

Answer by checkley79(3341)

P=2L+2W
L=3W-4
160=2(3W-4)+2W
160=6W-8+2W
8W=160+8
8W=168
W=168/8
W=21 cm.
L=3*21-4
L=63-4
L=59 cm.
PROOF:
160=2*59+2*21
160=118+42
160=160